OpenStudy (anonymous):

whats the way to solve this problem? l4x+1l=6

OpenStudy (anonymous):

http://www.twiddla.com/solved

OpenStudy (anonymous):

solve two equations: \[4x+1=6\] and \[4x+1=-6\] that is all

OpenStudy (anonymous):

14x=6-11 14x/14=-5/14 x=-5/14

jhonyy9 (jhonyy9):

4x+1=6 ------- 4x+1=-6 4x=5 4x=-7 x=5/4 x=-7/4